A metal cube(faces 20 mm+-0.1mm) fully sinks into a glass full of water. The diameter of the glass is 60 mm+-0,2 mm.Calculate the water rise level and its cumulative error.

1
Expert's answer
2022-01-16T17:56:03-0500

The water rise level:


"\Delta h=V_\text{cube}/A_\text{glass},\\\space\\
\Delta h=a^3/(\pi d^2/4)=\dfrac{4a^3}{\pi d^2}=2.8\text{ mm}."


The cumulative error: 0.1+0.2=0.3 mm.